# Representative Action Selection for Large Action Space Meta-Bandits

### v2

@@ 1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. Now, consider medicine demand: If a drug treats a very rare illness, it is unlikely to be needed frequently, so the store can exclude it to optimize storage space. Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
@@
Different from prior approaches in multi-armed bandits (MAB) (Dani et al., 2008) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
@@
Consider the following MAB setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (Vershynin, 2018, Chapter 7), and we also provide results under general sub-Gaussian assumption. Our underlying setting aligns with the one considered in contextual bandit (Dani et al., 2008); however, the key difference lies in the objective.
@@
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
@@
$\boxed{\textnormal{Regret}:=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}},}$ (1)
@@
This objective is motivated by practical considerations. The classic Bayesian regret in the bandit literature (Agrawal and Goyal, 2012) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower due to the reduced number of suboptimal actions. This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive. To see this, we can decompose the bandit regret as follows:
[…]
@@
In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (Agrawal and Goyal, 2012), in round $t$, and $N$ is the number of rounds. The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy. The inequality follows from the well-known regret bounds for Thompson Sampling (Lattimore and Szepesvári, 2020). $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant. Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
Our main contributions are organized as follows:
@@ 1.1 Related Works
Multi-Armed Bandits (Lai and Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal and Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
@@
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few independent samples as possible (Jamieson and Nowak, 2014; Kaufmann et al., 2016). Popular methods in fixed confidence setting include Action Elimination Even-Dar et al. (2006); Karnin et al. (2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive halving Karnin et al. (2013), successive reject Audibert and Bubeck (2010). Our algorithm includes a step for identifying the optimal action. While this line of work assumes (non-centered) sub-Gaussian outcome distributions within a bandit instance, we model each instance with a linear structure, allowing for further simplification.
@@
Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ Auer (2002); Dani et al. (2008); Rusmevichientong and Tsitsiklis (2010), and our model of each bandit instance align with this setting. This is a sequential method for maximizing a linear function—i.e., $f(a):=\mu_{a}(\theta)$ with fixed $\theta$—based on noisy observations of the function values. Soare et al. (2014) study the sample complexity of optimal action identification in this setting. However, this line of work assumes the action feature vectors are known.
@@
GP Optimization can deal with the case that the feature vectors of actions are unknown. Srinivas et al. (2009) models the outcome function as a sample from a GP prior with a kernel function (Williams and Rasmussen, 2006). Their Bayesian algorithm, GP-UCB, uses this GP prior to first select points $x_{t}$ that enable global estimation of the function, and then plays the point with the highest posterior mean. The widespread adoption of this method in bandit settings (Valko et al., 2013; Li and Scarlett, 2022), as well as in continuous action spaces (Chowdhury and Gopalan, 2017), highlights the practicality of assuming that action outcomes are correlated. This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
@@
Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight a advantage of our approach. These methods assume a fixed $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed. In contrast, our algorithm can adapt the subset size on the fly. For example, if 100 runs all return the same action, choosing this one may suffice ($|\mathcal{A}|=1$); if each run yields a different action, we may need more iterations to ensure a small regret ($|\mathcal{A}|>100$).
[…]
@@
So under the Best-222-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
@@
In contrast, our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$, which is the true optimal.
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB Chen et al. (2016), CTS Wang and Chen (2018). We argue that this line of work is not applicable to our framework: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
[…]
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability (Vershynin, 2018). The sharpest known bounds are due to Talagrand (2014).
@@ 2 Subset Selection Framework
$\mu_{a}:=\langle a,g\rangle+c,\quad\forall,a\in\mathcal{A}_{\operatorname*{full}},;g\sim\mathcal{N}(0,\Sigma),$ (2)
where $\Sigma$ is a positive semi-definite matrix. Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix. The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$. Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
@@
Therefore, the setting in Equation (2) is equivalent to
$\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$. Since the constant $c$ does not affect the regret (as defined in Equation (1)), we can, without loss of generality, focus on this canonical Gaussian process (Vershynin, 2018, Chapter 7) in the remainder:
@@
By Equation (3), it yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
@@ 2.1 Epsilon Nets
If the set of extreme points $\mathcal{A}_{\operatorname*{full}}$ is still large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions. This ensures that for every action in the full space, there exists a representative that is close to it. This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (Vershynin, 2018, Chapter 4) and another from a measure-theoretic perspective (Matousek, 2013, Chapter 10). Let $\|\cdot\|_{2}$ denote the Euclidean norm. Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
@@
A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small):
$\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$
where $\theta^{\top}$ denotes the transpose of $\theta$. The equalities use Equation (3) and $\mathbb{E}\theta\theta^{\top}=I$. Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance. However, this net suffers from the curse of dimension: e.g., for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.

### v3

@@ 1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. This correlation structure makes subset selection particularly interesting: if two drugs treat the same disease, choosing one may suffice, while drugs for distinct diseases might both be kept. In addition, demand also matters: drugs for rare illnesses can be excluded to save space, while flu medications should be stocked during flu season.
@@
Different from prior approaches in multi-armed bandits (MAB) (lai1985asymptotically, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
@@
Consider the following setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (vershynin2018high, , Chapter 7), while a more general sub-Gaussian assumption is considered in the Supplementary. Our setting aligns with that of contextual bandits (dani2008stochastic, ), and the restriction to linear functions may be partially mitigated by allowing $n$ to be unbounded.
@@
Consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$. For a given bandit instance $\theta$, if the optimal action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore. Conversely, if the optimal action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, for a bandit instance $\theta$, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
@@
$\boxed{\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),}$ (1)
@@
This objective is motivated by practical considerations. The classic Bayesian regret in the bandit literature (agrawal2012analysis, ) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower. This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive. To see this, we can decompose the bandit regret as follows:
$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
@@
In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (agrawal2012analysis, ), in round $t$, and $N$ is the number of rounds. The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy. The inequality follows from the well-known regret bounds for Thompson Sampling (lattimore2020bandit, ) [^1] This bound could be relaxed if the feature vector $a$ is known, or if $a$ can be incorporated as the input of a kernel function, as discussed in related works. However, we use this general bound because we also consider the most general case where $\mathcal{A}_{\operatorname*{full}}$ is simply an index set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$., where $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant. Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$. Our main contributions are organized as follows:
@@ 1.1 Related Works
Multi-Armed Bandits (lai1985asymptotically, ; auer2002finite, ) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. Popular methods include the UCB (auer2002finite, ), TS (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ) for adversarial settings.
@@
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (jamieson2014best, ; kaufmann2016complexity, ). Popular methods in fixed confidence setting include Action Elimination even2006action ; karnin2013almost , UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive halving karnin2013almost , successive reject audibert2010best .
@@
Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ auer2002using ; dani2008stochastic ; rusmevichientong2010linearly . This line of work assumes that the action feature vectors are known, so the cardinality of the action space does not play a role and may even be infinite. soare2014best study the sample complexity of optimal action identification in this setting, which scales linearly with the action space dimension $n$.
@@
GP Optimization addresses the case where the feature vectors of actions are unknown but a kernel function is available, so the inputs to the kernel are required. srinivas2009gaussian model the outcome function as a sample from a GP prior with a kernel function (williams2006gaussian, ), achieving bandit regret that scales as $\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}$ for finite action space. The widespread adoption of this method in bandit settings (valko2013finite, ; li2022gaussian, ), as well as in continuous action spaces (chowdhury2017kernelized, ), highlights the practicality of assuming that action outcomes are correlated. This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
@@
Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight an advantage of our approach. These methods assume a fixed subset cardinality $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed. In contrast, our algorithm can adapt the subset size on the fly.
Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (kalyanakrishnan2012pac, ; gabillon2012best, ; kaufmann2016complexity, ; chen2017nearly, ). This line of work assumes that all actions are independent and have distinct expected values, making its methods inapplicable to our framework. If one were to apply these methods regardless, the most reasonable approach, in our view, would be to treat the family of bandit instances as a super-bandit, where each bandit instance corresponds to a round, and the expected payoff of an action in that round is given by $\mu_{a}$. In this setting, top-$K$ identification would refer to selecting $K$ actions with the highest expected payoffs $\mathbb{E}[\mu_{a}]$. In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant—so that the entire action space shares the same highest expected payoff. Further, even if $\mathbb{E}[\mu_{a}]$ varies across actions, ignoring correlations can be fatal in our framework:
@@
So under the Best-$2$-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
@@
Our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$—the true optimal.
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB chen2016combinatorial , CTS wang2018thompson . We argue that this line of work is not directly applicable to our framework, but we include it as a baseline in subsequent empirical evaluations: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
@@ 2 Subset Selection Framework
$\mu_{a}:=\langle a,g\rangle+c,\quad\forall\;a\in\mathcal{A}_{\operatorname*{full}},\;g\sim\mathcal{N}(0,\Sigma),$ (2)
where $\Sigma$ is a positive semi-definite matrix. Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$. The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$. Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
@@
Therefore, the setting in equation 2 is equivalent to $\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$ where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$. Since the constant $c$ does not affect the regret (as defined in equation 1), we can, without loss of generality, focus on this canonical Gaussian process (vershynin2018high, , Chapter 7) in the remainder:
@@
By equation 3, it yields $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
@@ 2.1 Epsilon Nets
If $\mathcal{A}_{\operatorname*{full}}$ is very large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions. This ensures that for every action in the full space, there exists a representative that is close to it. This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (vershynin2018high, , Chapter 4) and another from a measure-theoretic perspective (matousek2013lectures, , Chapter 10). Let $\|\cdot\|_{2}$ denote the Euclidean norm. Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
@@
A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small): $\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$ where $\theta^{\top}$ denotes the transpose of $\theta$. The equalities use equation 3 and $\mathbb{E}\theta\theta^{\top}=I$. Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance. However, this net suffers from the curse of dimension: e.g., for $0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.